# BOSC: A Backdoor-based Framework for Open Set Synthetic Image Attribution

### v1

@@ VI-A GAN Face Editing Dataset and Setting
The dataset for these experiments is built as in [30]. The face images are taken from the CelebA-HQ dataset. The images are manipulated with 18 editing types: 4 facial attributes are edited with InterfaceGAN [53], and 14 facial attributes with StyleCLIP [54]. The reconstructed version of the image with no editing (’None’) by PTI [55] is also considered, for a total of 19 classes. An overview of the dataset is provided in Table V. The editing attributes are grouped into four categories: expression, ageing, hairstyle and identity change. The editing types are split into 11 in-set and 8 out-of-set classes. Three different splittings of in-set and out-of-set editing types are considered, named ’Config-F1’, ’Config-F2’ and ’Config-F3’. Fig. 9 shows an example of a manipulated face image for each edited attribute. A total number of 5992 images were taken from CelebA-HQ dataset. For each configuration, 4400 CelebA-HQ images have been used to generate the edited images used for training, 1592 for those used for testing, for a total of 83600 (4400 × 11) images for training and 30248 (1592 × 19) for testing.
@@ VI-B Results
TABLE VI: Performance of facial editing classification in closed-set (Accuracy) and open-set (AU-ROC, AU-OSCR). The best results are shown in bold (the second-best is underlined).
@@
Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR
ARPL [13] 91.92 86.84 82.54 94.41 87.34 84.57 90.99 85.84 80.64 92.44 86.67 82.58
AKPF [15] 94.41 91.09 87.84 95.33 88.72 86.49 91.45 87.35 82.58 93.73 89.05 85.64
PCSSR [14] 95.33 85.60 82.77 96.40 82.60 80.63 91.87 86.23 81.39 94.53 84.81 81.60
RCSSR [14] 95.05 82.46 79.48 97.02 89.98 88.37 93.27 82.68 78.41 95.11 85.04 82.09
Res50Vit [30] 93.65 91.42 87.65 95.59 91.66 89.50 91.83 86.64 82.13 93.69 89.91 86.43
Baseline 97.21 87.85 86.66 97.91 88.22 87.24 95.09 86.10 83.92 96.74 87.39 85.94
BOSC (Ours) 96.65 92.13 90.35 97.28 91.62 90.49 94.50 88.43 85.40 96.16 90.73 88.75
@@
The results achieved for in-set and out-of-set samples are shown in Table VI. We see that all methods obtain AU-ROC and AU-OSCR larger than 80%, while the closed-set accuracy of the various methods ranges between 92% and 97%, with BOSC performing the best in almost all the cases. Besides, the results of all methods are stable across the various configurations. The best open-set results for the state-of-the-art are achieved by AKPF and Res50Vit, with the latter yielding the best performance. Compared to Res50Vit, BOSC achieves a small gain of 0.82% in AU-ROC and 2.32% in AU-OSCR (and a 2.47% gain in the Accuracy). We stress that Res50Vit is specialized for this task and resorts to the aid of a localization branch to focus on the face regions that are most relevant for the various editing. Therefore, the similar (slightly better) performance obtained by BOSC represents a noticeable result, also proving its generality across different image forensic tasks.
@@ VII Conclusion
Future work will focus on investigating the impact of the choice of trigger images on the results and on optimizing the triggers for a given classification task. We also plan to apply BOSC to other open-set classification tasks, also beyond image forensics, to validate its effectiveness for general OSR image classification tasks.

### v2

@@ VI-A GAN Face Editing Dataset and Setting
The dataset for these experiments is built as in [35]. The face images are taken from the CelebA-HQ dataset. The images are manipulated with 18 editing types: 4 facial attributes are edited with InterfaceGAN [58], and 14 facial attributes with StyleCLIP [59], namely, methods for image manipulation based on StyleGAN that exploit the disentanglement of the features in the latent space, utilizing it to edit the semantic attributes through linear subspace projection. To get the desired editing, the Pivotal Tuning Inversion (PTI) method is considered to extract the latent code of the real images [60]. Then, the image attribute is manipulated via InterfaceGAN or StyleCLIP using the learned latent code and a target StyleGAN2 generator to achieve the desired modification of the attribute. An overview of the dataset is provided in Table VI. The reconstructed version of the image with no editing (‘None’) by PTI [60] is also considered, for a total of 19 classes. The editing attributes are grouped into four categories: expression, aging, hairstyle and identity change. The editing types are split into 11 in-set and 8 out-of-set classes. Three different splittings of in-set and out-of-set editing types are considered, named ‘Config-F1’, ‘Config-F2’ and ‘Config-F3’. Fig. 12 shows an example of a manipulated face image for each edited attribute. A total number of 5992 images were taken from CelebA-HQ dataset. For each configuration, 4400 CelebA-HQ images have been used to generate the edited images used for training, 1592 for those used for testing, for a total of 83600 (4400 × 11) images for training and 30248 (1592 × 19) for testing.
@@ VI-B Results
TABLE VII: Performance of facial editing classification in closed-set (Accuracy) and open-set (AU-ROC, AU-OSCR). The best results are shown in bold (the second-best is underlined). The average EER of out-of-set detection is also reported.
@@
Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC AU-OSCR Accuracy AU-ROC EER AU-OSCR
ARPL [16] 91.92 86.84 82.54 94.41 87.34 84.57 90.99 85.84 80.64 92.44 86.67 19.20 82.58
AKPF [18] 94.41 91.09 87.84 95.33 88.72 86.49 91.45 87.35 82.58 93.73 89.05 18.36 85.64
PCSSR [17] 95.33 85.60 82.77 96.40 82.60 80.63 91.87 86.23 81.39 94.53 84.81 21.31 81.60
RCSSR [17] 95.05 82.46 79.48 97.02 89.98 88.37 93.27 82.68 78.41 95.11 85.04 22.84 82.09
Res50Vit [35] 93.65 91.42 87.65 95.59 91.66 89.50 91.83 86.64 82.13 93.69 89.91 17.83 86.43
Baseline 97.21 87.85 86.66 97.91 88.22 87.24 95.09 86.10 83.92 96.74 87.39 18.03 85.94
BOSC (Ours) 96.65 92.13 90.35 97.28 91.62 90.49 94.50 88.43 85.40 96.16 90.73 16.49 88.75
@@
The results achieved for in-set and out-of-set classes are shown in Table VII. We see that all methods obtain AU-ROC and AU-OSCR larger than 80%, while the closed-set accuracy of the various methods ranges between 92% and 97%, with BOSC performing the best in almost all the cases. Besides, the results of all methods are stable across the various configurations. The best open-set results for the state-of-the-art are achieved by AKPF and Res50Vit, with the latter yielding the best performance. Compared to Res50Vit, BOSC achieves a small gain of 0.82% in AU-ROC, 1.34% in EER, and 2.32% in AU-OSCR (and a 2.47% gain in the Accuracy). We stress that, while AKPF (as well as ARPL and the CSSR methods) is a general algorithm for OSR, Res50Vit is a method designed for open-set facial editing classification, hence specialized for this task. This method resorts to the aid of a localization branch to focus on the face regions that are most relevant for the various editing. Therefore, the similar (slightly better) performance obtained by BOSC represents a noticeable result, also proving its generality across different image forensic tasks. In addition, we applied BOSC to this task using the same set of trigger images adopted for architecture attribution. However, better performance could be achieved with a different set (the optimization of the set of trigger images is left to future work).
With regard to the performance at a fixed threshold, BOSC achieves an average CCR of 93.6% and a TPR of 52.9% at FPR=5%. The threshold values $\nu$ are in the range 9-11, which is similar to the range of values for the architecture attribution task.
@@ VII Conclusion
Future work will focus on investigating the impact of the choice of trigger images on the results and on optimizing the triggers for a given classification task. Notably, our experiments revealed that not all triggers are equally effective in the context of OSIA. Unlike malevolent backdoor attacks, where stealthiness is a major requirement, OSIA triggers do not need to be hidden - making their visibility acceptable and even potentially beneficial. In fact, we observed that semantically meaningful triggers (e.g., cartoon images) consistently outperform low-level, texture-based triggers in rejecting unknown classes. This insight suggests that the nature of the trigger plays a crucial role in OSIA performance and opens a new research direction: the design of specific, semantically rich triggers to maximize open-set attribution performance. We also plan to apply BOSC to other open-set classification tasks, also beyond image forensics, to validate its effectiveness for general OSR image classification tasks.